After walking across a dry carpet, there's a potential difference of 4400 V between your hand and the doorknob. As you reach toward the knob, a spark jumps, releasing energy 1.0 μJ .
Answer #1

You only wrote the statement of the problem, but I found some questions that I think are what you want:

a.- How much charge was transferred?

The charge can be calculated with this expression for energy:

E = qV

Where q is charge and V is voltage. So:

q = \frac{E}{V}

q = \frac{1x10^{-6}J}{4400v}

ANSWER q = 2.27x10^{-10}C

b.- Assuming all the charges were electrons, how many were there?

Charge is given by:

q = ne

n = \frac{q}{e}

n = \frac{2.27x10^{-10}C}{1.602x10^{-19}C}

ANSWER n = 1.42x10^{9}electrons

I hope it helps!!